OpenStudy (anonymous):

An electron is accelerated from rest through a potential difference of 3750 V. It enters a region where B = 4.0 x 10^-3 T perpendicular to its velocity. Calculate the radius of the path it will follow. Would I use the equation: mv^2/r = qvB; where r = mv/qB ?

OpenStudy (souvik):

yes...be careful..here v is the velocity of the electron..not the potential difference..!

OpenStudy (anonymous):

hmm...ok... so m = mass of the electron, v = velocity of electron, q = charge of electron and B = magnetic force? is that right?

OpenStudy (anonymous):

and all are given besides V, so, do I need to find V first in order to use the first equation?

OpenStudy (souvik):

hm

OpenStudy (anonymous):

yep. im kinda stumped here ha.

OpenStudy (souvik):

the electron is accelerated from rest through a potential difference V (say) then the kinetic energy of the electron=work done 1/2mv^2=qV.....[v=velocity of electron , V=potential difference] mv^2=2qV (mv)^2=2qmV mv=(2qmV)^1/2 now ur eqn reduced to..r={(2qmV)^1/2}/qB

OpenStudy (anonymous):

hmm... ok I think I follow you on how you got to the final equation (eliminating velocity)... let me try to put in some numbers and see what I get. Thanks for the help

OpenStudy (souvik):

u r welcome..:)

OpenStudy (anonymous):

So, my final answer is 0.0516 ....meters? I think. Which would be 5.2 cm. Which is the answer I was looking for. I think this is correct. Again, thanks!!

OpenStudy (souvik):

:)